OpenStudy (anonymous):

James has 19 phone cards worth either $20 or $15. how many of each does she have if the phone cards total $300?

OpenStudy (anonymous):

300/19= 15

OpenStudy (anonymous):

You have to set up a system of equation. x+y=19 20x+15y=300 Solve for x and y.

OpenStudy (anonymous):

x+y=19 20x+15y=300 -20x-20y=-380 20x+15y=300 Add the equations together -5y=-80 y=16 Plug y=16 back into one of the equations x+y=19 x+16=19 x= 19-16 x=3 So she has 3 of the $20 cards and 16 or the $15 cards.

OpenStudy (anonymous):

ohh I see .. i was wrong then

OpenStudy (anonymous):

Clearly explained stargirlzilin, Thank You.